The heights of the population of boys are normally distributed with a mean of 66 inches and a 

standard deviation of 8.9 inches. If a random sample of 40 boys is drawn from this population,


1
Expert's answer
2022-05-10T13:09:37-0400
"P(X>66)=1-P(Z\le\dfrac{66-66}{8.9/\sqrt{40}})"

"=1-P(Z\le 0)=0.5"

"P(Z>64.5)=1-P(Z\le\dfrac{64.5-66}{8.9/\sqrt{40}})"

"\approx1-P(Z\le-1.065936)\approx0.8568"



"P(64.5<X<66)=P(Z<\dfrac{66-66}{8.9/\sqrt{40}})"

"-P(Z\le\dfrac{64.5-66}{8.9/\sqrt{40}})\approx0.5-0.143226"

"\approx0.3568"